Find dy/dx by implicit differentiation

tan(x-y)=y/1+x^2


1
Expert's answer
2022-04-16T04:13:04-0400

"\tan(x-y)=\frac{y}{1+x^2}"

"(\tan(x-y))’=(\frac{y}{1+x^2})’"

"\frac{(x-y)’}{\cos^2(x-y)}=\frac{y’(1+x^2)-y(1+x^2)’}{(1+x^2)^2}"

"\frac{-y’}{\cos^2(x-y)}=\frac{y’(1+x^2)-2xy}{(1+x^2)^2}"

"-y’(1+x^2)^2=(y’(1+x^2)-2xy) \cos^2(x-y)"

"y’=\frac{2xy\cos^2(x-y)}{(1+x^2)(\cos^2(x-y)+(1+x^2))}"